OpenStudy (anonymous):

hmmm.. excuse me everyone. may i ask a question? 1/2 = 2 prove please qive me an answer tnx ..

jhonyy9 (jhonyy9):

\[2^{-1} = 2^{1} ??? \]

jhonyy9 (jhonyy9):

you wann to prove this ???

jhonyy9 (jhonyy9):

if 1/2 =2 so than 1/2=4/2 so than 1=4 so this not is right ,not is equal

jhonyy9 (jhonyy9):

hope that this is understandably

OpenStudy (anonymous):

1/2 = 2 ------1/2 = 50% ------- 2/1 = 2 (Bookmakers Percentages & Price for Gambling) Could be the correct option.

OpenStudy (anonymous):

i don't know, but it's going to be a weird proof for sure.

OpenStudy (amistre64):

when i cut a pie in half, i have 2 pieces